OpenStudy (anonymous):

help solving using substitution?? 2x+4y=3 and 4x+4y=8

OpenStudy (anonymous):

4x+4y=8 divide both sides by 4 x+y=2 x=2-y

OpenStudy (anonymous):

substitute it to 2x+4y=3 2(2-y)+4y=3 4-2y+4y=3 2y=-1 u can find y, divide both sides by 2 to find x, substitute y u have got to x+y=2 happy to help :)